# Generalization Error of the Tilted Empirical Risk

### v2

@@ 3.1 Uniform Bounds
Let $A(\gamma)=(1-\exp(\gamma M))^{2}$. Under the same assumptions in Theorem 3.3, with probability at least $(1-\delta)$, and a finite hypothesis space, the absolute value of the titled generalization error satisfies
$\begin{split}&\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|\\ &\leq\frac{\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}\\ &\qquad+\frac{\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
where $A(\gamma)=(1-\exp(\gamma M))^{2}$.
Corollary 3.7.
Under the assumptions in Corollary 3.6, if $\gamma$ is of order $O(n^{-\beta})$ for $\beta>0$, then, as $A(\gamma)\sim\gamma^{2}M^{2}$ for $\gamma\rightarrow 0$, the upper bound on the absolute tilted generalization error in Corollary 3.6 has a convergence rate of $\max\big{(}O(1/\sqrt{n}),O(n^{-\beta})\big{)}$ as $n\rightarrow\infty$.
Remark 3.8.
Choosing $\beta\geq 1/2$ in Corollary 3.7 gives a convergence rate of $O(1/\sqrt{n})$ for the tilted generalization error.
Remark 3.9 (The influence of $\gamma$).
As $\gamma\rightarrow 0$, the upper bound in Corollary 3.6 on the absolute value of tilted generalization error converges to the upper bound on absolute value of the generalization error under the ERM algorithm obtained by Shalev-Shwartz and Ben-David (2014),
In particular, ${\big{(}\exp(|\gamma|M)-1\big{)}}/{|\gamma|}\rightarrow M$ and the first term in Corollary 3.6 vanishes. Therefore, the upper bound converges to a uniform bound on the linear empirical risk.
Remark 3.10.
The term $-\exp(-2\gamma M)/\gamma$ in Theorem 3.3 can cause the upper bounds in Theorem 3.3 to be tighter for $\gamma>0$ than for $\gamma<0$. Furthermore, in comparison with the uniform upper bound on the linear generalization error, (15), we obtain a tilted generalization error upper bounds which can be tighter, for sufficiently large values of samples $n$ and small values of $\gamma$.
Using Corollary 3.7, we derive an upper bound on the excess risk.
Corollary 3.11.
Under the same assumptions in Theorem 3.3, and a finite hypothesis space, with probability at least $(1-\delta)$, the excess risk of tilted empirical risk satisfies
$\begin{split}\mathfrak{E}_{\gamma}(\mu)&\leq\frac{2\big{(}\exp(|\gamma|M)-1\big{)}}{|\gamma|}\sqrt{\frac{\log(\mathrm{card}(\mathcal{H}))+\log(2/\delta)}{2n}}\\ &\qquad+\frac{2\max(1,\exp(-2\gamma M))A(\gamma)}{8|\gamma|},\end{split}$
where $A(\gamma)=(1-\exp(\gamma M))^{2}$.
The theorems in this section assumed that the hypothesis space is finite; this is for example the case in classification problems with a finite number of classes. If this assumption is violated, we can apply the growth function technique from (Bousquet et al., 2003; Vapnik, 1999). Furthermore, the growth function can be bounded by VC-dimension in binary classification (Vapnik, 1999) or Natarajan dimension (Holden and Niranjan, 1995) for multi-class classification scenarios. Note that the VC-dimension and Rademacher complexity bounds are uniform bounds and are independent of the learning algorithms.
3.2 Information-theoretic Bounds
Next, we provide an upper bound on the expected tilted generalization error. All proof details are deferred to Appendix D.2. For information-theoretic bounds, we employ the following decomposition of the expected tilted generalization error,
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\\ &=\{\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\}\\ &+\{\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\}.\end{split}$ (16)
The following is helpful in deriving the upper bound.
Proposition 3.12.
Under Assumption 3.1, the following inequality holds for any learning algorithm, $P_{H|S}$,
$\begin{split}&\Big{|}\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\Big{|}\\ &\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}.\end{split}$ (17)
Using Proposition 3.12, we derive the following upper and lower bounds on the expected generalization error.
Theorem 3.13.
Under Assumption 3.1, the expected tilted generalization error satisfies
$\overline{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}-\frac{\gamma\exp(-\gamma M)}{2}\Big{(}1-\frac{1}{n}\Big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.$ (18)
Theorem 3.14.
Under the same assumptions in Theorem 3.13, the expected tilted generalization error satisfies
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)\geq-\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}\\ &\qquad-\frac{\gamma\exp(\gamma M)}{2}\Big{(}1-\frac{1}{n}\Big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$
Combining Theorem 3.13 and Theorem 3.14, we derive an upper bound on the absolute value of the expected tilted generalization error.
Corollary 3.15.
Under the same assumptions in Theorem 3.13, the absolute value of the expected titled generalization error satisfies
$\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}\\ &\qquad+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
Remark 3.16.
In Corollary 3.15, we observe that by choosing $\gamma=O(n^{-\beta})$, the overall convergence rate of the generalization error upper bound is $\max(O(1/\sqrt{n}),O(n^{-\beta}))$ for bounded $I(H;S)$. For $\beta\geq 1/2$, the convergence rate of (18) is the same as the convergence rate of the expected upper bound in (Xu and Raginsky, 2017). In addition, for $\gamma\rightarrow 0$, the upper bound in Corollary 3.15 converges to the expected upper bound in (Xu and Raginsky, 2017).
The results in this section are non-vacuous for bounded $I(H;S)$. If this assumption is violated, we can apply the individual sample method (Bu et al., 2020), chaining methods (Asadi et al., 2018), or conditional mutual information frameworks (Steinke and Zakynthinou, 2020) to derive tighter upper bound for the tilted generalization error.
4 Generalization Bounds for Unbounded Loss Functions
In the previous section, we assumed that the loss function is bounded which is limiting in practice. Several works already proposed some solutions to overcome the boundedness assumption under linear empirical risk (Lugosi and Neu, 2022; Haddouche and Guedj, 2022; Alquier and Guedj, 2018; Holland, 2019) via a PAC-Bayesian approach. In this section, we derive upper bounds on the tilted generalization error via uniform approach for negative tilt ($\gamma<0$) under bounded second moment assumption with convergence rate of $O(1/\sqrt{n})$. In particular, we relax the bounded assumption (Assumption 3.1). The following assumptions are made for the uniform analysis.
Assumption 4.1 (Uniform bounded second moment).
There is a constant $\kappa_{u}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\mu}[\ell^{2}(h,Z)]\leq\kappa_{u}^{2}$ uniformly for all $h\in\mathcal{H}$.
The assumption on second moment, Assumption 4.1, can be satisfied if the loss function is sub-Gaussian or sub-Exponential (Boucheron et al., 2013) under the distribution $\mu$ for all $h\in\mathcal{H}$. All proof details for the results in this section are deferred to Appendix E. In this following we present the results based on uniform bounds. The results for unbounded loss function based on information theoretic bounds are presented in Appendix E.2.
For unbounded loss function, we consider the decomposition of the tilted generalization error in (12) for $\gamma<0$. The term $I_{1}$ can be bounded using Lemma C.9. Then, we apply Bernstein’s inequality (Boucheron et al., 2013) to provide upper and lower bounds on the second term, $I_{2}$.
Theorem 4.2.
Given Assumption 4.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$,
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq 2\kappa_{u}\exp(-\gamma\kappa_{u})\sqrt{\frac{\log(2/\delta)}{n}}\\ &-\frac{4\exp(-\gamma\kappa_{u})\log(2/\delta)}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (19)
Theorem 4.3.
Given Assumption 4.1, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, and $\gamma<0$, the following lower bound on the tilted generalization error holds
$\begin{split}\mathrm{gen}_{\gamma}(h,S)&\geq-\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}\\ &+\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}.\end{split}$ (20)
Combining Theorem 4.2 and Theorem 4.3, we derive an upper bound on the absolute value of the tilted generalization error.
Corollary 4.4.
Under the same assumptions in Theorem 4.3 and a finite hypothesis space, then for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies
$\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|\leq\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{B(\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{u})B(\delta)}{3n\gamma(1-\zeta)}-\frac{\gamma}{2}\kappa_{u}^{2},$ (21)
@@ 4 Generalization Bounds for Unbounded Loss Functions
For $\gamma\asymp n^{-1/2}$, $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(-2\kappa_{u})}$ and $-1<\gamma<0$, the upper bound in Corollary 4.4 gives a theoretical guarantee on the convergence rate of $O(n^{-1/2})$. Using TER with negative $\gamma$ can help to derive an upper bound on the absolute value of the tilted generalization error under the bounded second-moment assumption.
Similar to Corollary 3.11, an upper bound on excess risk under unbounded loss function assumption can be derived.
Corollary 4.6.

### v3

@@ 3.1 Uniform Bounds
Under the same assumptions in Proposition 3.4 and a finite hypothesis space, then for $n\geq\frac{(4|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$, for $\gamma<0$ and with probability at least $(1-\delta)$, the absolute value of the titled generalization error satisfies
$\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|$ (15)
$\leq$ $\frac{2\exp(|\gamma|\kappa_{u})}{(1-\zeta)|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{u}^{1+\epsilon}B(\delta)}{n}}+\frac{4\exp(|\gamma|\kappa_{u})B(\delta)}{3n|\gamma|(1-\zeta)}$
$+|\gamma|^{\epsilon}\kappa_{u}^{1+\epsilon},$
@@
For $\gamma\asymp n^{-1/(1+\epsilon)}$, $n\geq\frac{(4\kappa_{u}^{1+\epsilon}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(-2\kappa_{u})}$ and $-1<\gamma<0$, the upper bound in Theorem 3.5 gives a theoretical guarantee on the convergence rate of $O(n^{-\epsilon/(1+\epsilon)})$. Using TER with negative tilt can help to derive an upper bound on the absolute value of the tilted generalization error under the bounded $(1+\epsilon)$-th moment assumption.
The following Lemma establishes an upper bound on the excess risk of tilted empirical risk, expressed in terms of $\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|$.
Lemma 3.7.
The excess risk of the tilted empirical risk satisfies,
$\mathfrak{E}_{\gamma}(\mu)\leq 2\sup_{h\in\mathcal{H}}|\mathrm{gen}_{\gamma}(h,S)|.$ (16)
Combining Lemma 3.7 with Theorem 3.5, an upper bound on the excess risk of the tilted empirical risk can be derived (See Appendix D.1).
The theorems in this section assume that the hypothesis space is finite; this is, for example, the case in classification problems with a finite number of classes. If this assumption is violated, we can apply the growth function technique from Bousquet et al. (2003); Vapnik (1999). In particular, the growth function can be bounded by VC-dimension in binary classification (Vapnik, 1999) or Natarajan dimension (Holden & Niranjan, 1995) for multi-class classification scenarios. Note that the VC-dimension and Rademacher complexity bounds are uniform bounds and are independent of the learning algorithms. Furthermore, one can construct an $\epsilon$-net over the hypothesis space, thereby discretizing the space. In this construction, we select a finite subset $H^{\prime}\subset\mathbb{R}^{m}$ such that for every $h\in H$, there exists a $h^{\prime}\in H^{\prime}$ with $\|h-h^{\prime}\|\leq r$. By applying our finite hypothesis result to this discretized set and controlling the approximation error through the Lipschitz property of the loss function, we effectively generalize our bounds to the continuous case. This method is exemplified in (Xu & Raginsky, 2017), which demonstrates that “for an uncountable hypothesis space, we can always convert it to a finite one by quantizing the output of the smallest set $H^{\prime}$ such that for all $h\in H$ there is a $h^{\prime}\in H^{\prime}$ with $\|h-h^{\prime}\|\leq r$, where the Lipschitz maximal inequality (Lemma 5.7 in (Vershynin, 2010)) is derived using a similar quantization technique.
3.2 Information-theoretic Bounds
Assumption 3.8 (Bounded $(1+\epsilon)$-th moment).
The learning algorithm $P_{H|S}$, loss function $\ell$, and $\mu$ are such that there is a constant $\kappa_{t}\in\mathbb{R}^{+}$ with which the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\max\big{(}\mathbb{E}_{P_{H,Z}}[\ell^{\epsilon}(H,Z)],\mathbb{E}_{P_{H}\otimes\mu}[\ell^{1+\epsilon}(H,Z)]\big{)}\leq\kappa_{t}^{1+\epsilon}$ for some $\epsilon\in(0,1]$.
For our information-theoretical analysis, we use the following decomposition;
$\begin{split}&\overline{\mathrm{gen}}_{\gamma}(H,S)=\underbrace{\mathrm{R}(H,P_{H}\otimes\mu)-\mathrm{R}_{\gamma}(H,P_{H}\otimes\mu)}_{I_{3}}\\ &\quad+\widehat{\mathrm{gen}}_{\gamma}(H,S)\\ &\quad+\underbrace{\mathrm{R}_{\gamma}(H,P_{H,S})-\mathbb{E}_{P_{H,S}}[\widehat{\mathrm{R}}_{\gamma}(H,S)]}_{I_{4}}.\end{split}$ (17)
We can derive bounds on $I_{3}$ and $I_{4}$ in a similar approach to Proposition 3.2.
Then, we provide upper and lower bounds on the non-linear expected generalization error of the tilted empirical risk, $\widehat{\mathrm{gen}}_{\gamma}(H,S)$, under negative tilt. These results are helpful in deriving an upper bound on the absolute expected tilted generalization error.
Proposition 3.9.
Given Assumption 3.8 and assuming that $\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}\leq n$ holds, then the following inequality holds,
$\begin{split}\widehat{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\sqrt{\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}},\end{split}$
and if $\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}>n$ holds, then following inequality holds,
$\begin{split}\widehat{\mathrm{gen}}_{\gamma}(H,S)\leq\frac{\exp(|\gamma|\kappa_{t})}{|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}.\end{split}$
Proposition 3.10.
Under Assumption 3.8 for some $\epsilon\in(0,1]$, assume further that there exists $\zeta\in(0,1)$ such that one of the following cases holds,
(a) $\zeta\leq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{\zeta^{2}\exp(2\gamma\kappa_{t})}$,
(b) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and $n\geq\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}$,
(c) $\zeta>\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}\exp(|\gamma|\kappa_{t})}{2}$ and
$\frac{2I(H;S)}{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}>n\geq\frac{2I(H;S)}{2\zeta\exp(\gamma\kappa_{t})-2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}.$
Then, the following lower bounds on non-linear expected generalization error hold,
$\begin{split}\widehat{\mathrm{gen}}_{\gamma}(H,S)\geq\begin{cases}\frac{-\exp(|\gamma|\kappa_{t})}{|\gamma|(1-\zeta)}\sqrt{\frac{2|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}I(H;S)}{n}},\\ \quad\quad\mbox{ \text{if (a) or (b) hold,}}\\ \frac{-\exp(|\gamma|\kappa_{t})}{(1-\zeta)|\gamma|}\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)},\\ \quad\quad\mbox{ \text{if (c) holds}}.\end{cases}\end{split}$
Using Proposition 3.9 and Proposition 3.10, we can obtain upper bound and lower bounds on the expected tilted generalization error, respectively. Then, combining upper and lower bounds on expected tilted generalization error, we can derive the upper bound on the absolute value of the expected tilted generalization error in the following.
Theorem 3.11.
Under the same assumptions as in Proposition 3.10, the following upper bounds on absolute expected tilted generalization error hold,
$\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|\leq\begin{cases}D(\gamma)\sqrt{\frac{2\kappa_{t}^{\epsilon+1}|\gamma|^{1+\epsilon}I(H;S)}{n}}+C(\gamma),\\ \quad\quad\mbox{ \text{if (a) or (b) hold,}}\\ D(\gamma)\Big{(}\frac{I(H;S)}{n}+\frac{2^{\epsilon}|\gamma|^{1+\epsilon}\kappa_{t}^{1+\epsilon}}{2}\Big{)}+C(\gamma),\\ \quad\quad\mbox{ \text{if (c) holds}}.\end{cases}\end{split}$
where $C(\gamma)=|\gamma|^{\epsilon}\kappa_{t}^{1+\epsilon}$ and $D(\gamma)=\frac{\exp(|\gamma|\kappa_{t})}{|\gamma|(1-\zeta)}$.
Convergence rate: Assuming $\gamma=O(n^{-1/(1+\epsilon)})$ and $n\geq\frac{\kappa_{t}^{1+\epsilon}\exp(2\kappa_{t})I(H;S)}{\zeta^{2}}$, then the upper bound in Theorem 3.11 has the convergence rate $O(n^{-\epsilon/(1+\epsilon)})$. Note that the result in Theorem 3.11 holds for unbounded loss functions, provided that the $(1+\epsilon)$-th moment of the loss function exists for some $\epsilon\in(0,1]$.
The results in this section are non-vacuous for bounded $I(H;S)$. If this assumption is violated, we can apply the individual sample method (Bu et al., 2020), chaining methods (Asadi et al., 2018), or conditional mutual information frameworks (Steinke & Zakynthinou, 2020) to derive a tighter upper bound for the expected tilted generalization error.
4 Robustness of TER
Previous experimental work by Li et al. (2021, 2023a) has demonstrated that tilted empirical risk with negative tilt ($\gamma<0$) exhibits robustness against noisy samples and outliers during training. In this section, we investigate the robustness of TER under distributional shift scenarios. Inspired by the concept of influence functions (Marceau & Rioux, 2001; Christmann & Steinwart, 2004; Ronchetti & Huber, 2009), we model distributional shift through a distribution $\tilde{\mu}\in\mathcal{P}(\mathcal{Z})$, which represents the effect of outliers or noise in the noisy training dataset $\hat{S}$.
Uniform Approach: Our robustness analysis of TER is based on the following assumption.
Assumption 4.1 (Uniform bounded second moment under $\tilde{\mu}$).
There is a constant $\kappa_{s}\in\mathbb{R}^{+}$ such that the loss function $(H,Z)\mapsto\ell(H,Z)$ satisfies $\mathbb{E}_{\tilde{\mu}}[\ell^{1+\epsilon}(h,Z)]\leq\kappa_{s}^{1+\epsilon}$ uniformly for all $h\in\mathcal{H}$.
For our robustness analysis, we employ the following decomposition of the tilted generalization error,
$\begin{split}\mathrm{gen}_{\gamma}(h,\hat{S})&=\underbrace{\mathrm{R}(h,\mu)-\mathrm{R}_{\gamma}(h,\mu)}_{I_{1}}+\widehat{\mathrm{gen}}_{\gamma}(h,\hat{S})\\ &+\underbrace{\mathrm{R}_{\gamma}(h,\mu)-\mathrm{R}_{\gamma}(h,\tilde{\mu})}_{I_{4}}.\end{split}$ (18)
Using the functional derivative (Cardaliaguet et al., 2019), see also Appendix C, we can bound $I_{4}$ as follows.
Proposition 4.2.
Given Assumption 4.1 and Assumption 3.1, then the difference of tilted population risk under, (9), between $\mu$ and $\tilde{\mu}$ is bounded as follows for all $h\in\mathcal{H}$,
$\frac{1}{|\gamma|}\Big{|}\log(\mathbb{E}_{Z\sim\mu}[C(h,Z)])-\log(\mathbb{E}_{\tilde{Z}\sim\tilde{\mu}}[C(h,\tilde{Z})])\Big{|}$ (19)
$\leq\frac{\mathbb{TV}(\mu,\tilde{\mu})}{\gamma^{2}}\frac{\Big{(}\exp(|\gamma|\kappa_{u})-\exp(|\gamma|\kappa_{s})\Big{)}}{(\kappa_{u}-\kappa_{s})},$
where $C(h,Z)=\exp(\gamma\ell(h,Z))$.
Using Proposition 4.2, we can provide upper and lower bounds on the tilted generalization error under distributional shift. Then, combining upper and lower bounds, we can derive an upper bound on the absolute value of the tilted generalization error under distribution shift.
Theorem 4.3.